# Noise Contrastive Estimation-based Matching Framework for Low-resource Security Attack Pattern Recognition

### v1

@@ 5.2 Metrics and Baselines
The following baselines are targeted: Okapi BM25, adjusted from Husari et al. (2017). The bag-of-words model is enhanced by a security GloVe LM and TTP textual profile are used.
@@ 5.5 Ablation Studies
Hierarchical Labeling. We analyze the contributions of our hierarchical modeling to the ranking performances. As shown in Table 2, in general, our joint learning with the auxiliary task gives a notable performance boost in most scenarios. We report further in Table 3 the models’ results in the technique-level of the label hierarchy, where a sub-technique label is resolved to its technique. This is also a common practice in literature to streamline the complexity of the task. Overall, all models present significant improvements in this setting. Interestingly, the $\alpha$-balanced model, without the auxiliary task, is the best performer on the Expert dataset, under the technique-level setting. This is, nonetheless, understandable as the original hierarchical structure is semantically one level reduced in this case.
Transformers. We observe the positive contributions of SecBERT to the performance of all models in most cases. Nevertheless, without SecBERT (i.e., ()(-)( - ) Transformers), our models still very much on par with the strong XMTC baselines at $k=1$ and outperform them at $k=3$, indicating the better ranking capability, specially on the Expert dataset.
@@
Loss Analysis. In Figure 4, we present additional analysis on the impact of the size of negative samples. The results indicate that as the size increases, the model tends to converge faster and exhibit better performance. However, it appears that there are no additional benefits beyond a size of 60, which corresponds to 10% of the label space.
A further analysis on the score distribution of the ranked lists are reported in Figure 5. The details are provided in the caption for convenient reference.
@@
Table 7: Technique-level results on the entire Expert dataset, trained on the training splits of Procedure+ and Tram. Bold denotes best-performer.
Methods P@1 R@1 F1@3 MMR@3 F1@5 MRR@5
@@ Appendix A The Task of TTP Mapping
Although others exist, MITRE ATT&CK999 https://attack.mitre.org/  Storm et al. (2020) is the prevalent knowledge base and taxonomy of TTPs used in the literature. It currently comprises 14 tactics, 196 techniques, 411 subtechniques and thousands of procedures, continually curated by community experts.
@@ Appendix B Convergence Analysis
Based on the stability of the NCE losses, we briefly discuss the convergence properties of our adjusted losses.
@@
Boundedness of Gradients. Proof: Let $g(x,y)$ be the matching function such that $0\leq g(x,y)\leq 1$ for all $(x,y)$. Consider the NCE loss, i.e., @@@@-balanced with a scaling factor $\gamma$:
We want to prove that the gradients of the NCE loss with respect to the model parameters are bounded. Let $\nabla J_{\text{NCE}}(\theta)$ denote the gradient vector. Taking the partial derivative of $J_{\text{NCE}}(\theta)$ with respect to a parameter $\theta_{i}$, we have:
Using the linearity of the derivative, we can rewrite the above expression as:
@@
$\frac{\partial J_{\text{NCE}}(\theta)}{\partial\theta_{i}}=E_{p(x,y)}\left[\frac{\partial}{\partial\theta_{i}}\log g(x,y)\right]$
$-\gamma E_{p(x)}\left[\frac{\partial}{\partial\theta_{i}}\log\sum_{j}g(x,y_{j})\right]$
Since $0\leq g(x,y)\leq 1$, the derivative of $\log g(x,y)$ with respect to any parameter $\theta_{i}$ is bounded between 0 and 1. Similarly, the derivative of $\log\sum_{j}g(x,y_{j})$ with respect to $\theta_{i}$ can be bounded by considering the partial derivatives of $g(x,y_{j})$.
Therefore, we can conclude that:
$\left|\frac{\partial J_{\text{NCE}}(\theta)}{\partial\theta_{i}}\right|\leq\max\{1,\gamma\max_{x,y_{j}}|\partial_{\theta_{i}}g(x,y_{j})|\}$
The above inequality implies that the absolute value of the partial derivative of the NCE loss with respect to any model parameter is bounded by a finite value, scaled by $\gamma$. Hence, we have shown that the gradients of @@@@-balanced with the scaling factor $\gamma$ are bounded. The proof for the asymmetric loss can be derived in an analogous manner.
Lemma 1
The matching function $g(z_{i},z_{j})$ is Lipschitz-continuous with a constant $C$, meaning that for any $z_{i},z_{i}^{\prime},z_{j}$, we have $|g(z_{i},z_{j})-g(z_{i}^{\prime},z_{j})|\leq C|z_{i}-z_{i}^{\prime}|$.
Informal proof. Our Siamese neural networks-based matching function $g(z_{i},z_{j})\in[0,1]$. $\square$.
Lemma 2
The noise distribution $q$ satisfies the matching moment condition of the true distribution $p$, which, in essence, indicates that the covariance matrices of the two are similar.
Informal proof. Since the noise distribution is sampled over the whole corpus, the lemma holds true for the random sampling strategy. $\square$.
Thus, our loss is also Lipschitz-continuous and retains convergence properties of the original NCE losses, when optimized using SGD together with the random negative sampling.
@@ D.2 Training Procedure and Hyperparameters
D.3 Analysis
D.4 The Large Language Model (LLM) Dilemma.
In this section, our aim is to explore the extent to which information extraction tasks, particularly TTP mapping, can be addressed by LLMs, e.g., GPT-4 or PaLM-2.
@@ D.3 Analysis
In the pursuit of this target, we carefully crafted a setting to prompt ChatGPT (public version) for the a handful of samples in the Expert dataset. In general, the responses provided by Chat-GPT are remarkable and somewhat persuasive in certain instances. However, it is evident that the answers primarily consist of high-level information (sometimes hallucinatory), with a lack of granularity that makes it useful, e.g., for accurate modeling of the attack steps. Our objectives are, on the other hand, to precisely map to (sub- ) techniques so that to reveal the actual capabilities of a threat group or a particular attack.
Our findings are:
  • Chat-GPT: generate an answer that derives from the problem description, so-called prompt, which is rather non-deterministic, with no fixed answer for the same problem. This is partially due to the non-constrained answer search space.
  • What we need: generate best-matched TTPs by comparing the given text to the ATT&CK TTP collection (bounded, deterministic). However, LLMs in general have the potential to alleviate this issue to some extent if they undergo effective domain-specific training during the pre-training phase and are subsequently fine-tuned using high-quality datasets, such as an enhanced version of our Expert, in the reinforcement-learning-from-human-feedback phase Karpathy (2023).
  • Currently, Chat-GPT (3.5) is not up to the task.
Here are a couple of examples:
@@
inline,color=green!20!whiteinline,color=green!20!whitetodo: inline,color=green!20!whiteQ1: What MITRE ATT&CK techniques (TTPs) are mentioned in the following text: “After the .NET PE file has been run, we observed the same behavior as the above QUADAGENT sample of dropping a PowerShell script with the filename SystemDiskClean.ps1 alongside a VBScript file with the same name. The C2 techniques remained identical, with the only change being the server which became cpuproc[.]com.inline,color=purple!20!whiteinline,color=purple!20!whitetodo: inline,color=purple!20!whiteA1: The text mentions the following MITRE ATT&CK techniques: File and Directory Discovery: The QUADAGENT sample drops a PowerShell script and a VBScript file with the filename SystemDiskClean.ps1. Command and Control (C2) Communications: The text indicates that there is a change in the C2 server from appexsrv[.]net to cpuproc[.]com. It’s important to note that there may be additional MITRE ATT&CK techniques employed in this attack, but they are not mentioned in this particular text. A more detailed analysis would be required to confirm the full set of techniques used in this attack.
Implications: The answer from ChatGPT is rather generic, while missing out the compelling Masquerading technique.
inline,color=green!20!whiteinline,color=green!20!whitetodo: inline,color=green!20!whiteQ2: What MITRE ATT&CK techniques (TTPs) are mentioned in the following text: “Russian state-sponsored APT actors have exfiltrated credentials and exported copies of the Active Directory database ntds.dit.”inline,color=purple!20!whiteinline,color=purple!20!whitetodo: inline,color=purple!20!whiteA2: The following MITRE ATT&CK TTPs are mentioned in the given text: Exfiltration: The adversaries have exfiltrated credentials and data from the victim’s network. Collection: The adversaries have exported copies of the Active Directory database ntds.dit, which is a technique used for data collection.
Implications: Chat-GPT gives a rather generic answer, and cannot point to specific TTPs (i.e., NTDS or OS Credential Dumping).

### v4

@@ 5.2 Metrics and Baselines
The following baselines are targeted: Okapi BM25, adjusted from Husari et al. (2017). The BoW is augmented with $k$ closest terms from a security GloVe model, enhancing the BM25 retrieval capability. Here, query represents the target text, and documents refer to TTP descriptions.
@@ 5.5 Ablation Studies
Hierarchical Labeling. We analyze the contributions of our hierarchical modeling to the ranking performances. As shown in Table 2, in general, our joint learning with the auxiliary task gives a notable performance boost in most scenarios. We report further in Table 3 the models’ results in the technique-level of the label hierarchy, where a sub-technique label is resolved to its technique. This is also a common practice in literature to streamline the complexity of the task. Overall, all models present significant improvements in this setting. Interestingly, here the $\alpha$-balanced model, without the auxiliary task, is the best performer on the Expert dataset. This is, nonetheless, understandable as the original hierarchical structure is semantically one level reduced in this case.
Transformers. We observe the positive contributions of SecBERT to the performance of all models in most cases. Nevertheless, without SecBERT (i.e., w/o Transformers), our models are still very much on par with the strong XMTC baselines at $k=1$ and outperform them at $k=3$, indicating the better ranking capability, specially on the Expert dataset.
@@
Loss Analysis. In Fig. 4, we present additional analysis on the impact of the size of negative samples. The results indicate that as the size increases, the model tends to converge faster and exhibit better performance. However, it appears that there are no additional benefits beyond a size of 60, which corresponds to 10% of the label space.
A further analysis on the score distribution of the ranked lists are reported in Fig. 5. The details are provided in the caption for convenient reference.
@@
Table 7: Technique-level results on the entire Expert dataset. Legend in Table 6 applies.
Methods P@@@@1 R@@@@1 F1@@@@3 MMR@@@@3 F1@@@@5 MRR@@@@5
@@ Appendix A The Task of TTP Mapping
Although others exist, MITRE ATT&CK999 https://attack.mitre.org/  Storm et al. (2020) is the prevalent knowledge base and taxonomy of TTPs used in the literature. The version 12.0 comprises 14 tactics, 196 techniques, 411 sub-techniques and thousands of procedures, continually curated by community experts.
@@ Appendix B Convergence Analysis
We analyze the convergence properties of our NCE losses by establishing boundedness of gradients and Lipschitz continuity, crucial for SGD-based optimization.
@@
Boundedness of Gradients. Proof Sketch: Let $g_{\theta}(x,y)$ be the matching function. For the $\alpha$-balanced NCE loss:
$\mathsf{J}_{NCE}^{global}(\theta)=-\mathsf{E}_{(x,y)}\Big{[}$ $g_{\theta}(x,y)-\log\gamma$
$-\log\left(\sum_{j}\exp(g_{\theta}(x,y_{j}))\right)\Big{]}$
The gradient with respect to parameter $\theta_{k}$ is:
$\frac{\partial\mathsf{J}_{NCE}^{global}}{\partial\theta_{k}}=-\mathsf{E}_{(x,y)}\Big{[}$ $\frac{\partial g_{p}}{\partial\theta_{k}}$
$-\sum_{j}\frac{\exp(g_{n,j})}{\sum_{m}\exp(g_{n,m})}\frac{\partial g_{n,j}}{\partial\theta_{k}}\Big{]}$
Assuming $|\partial g_{\theta}(x,y)/\partial\theta_{k}|\leq B_{g}$ for some finite $B_{g}$, and noting that $\sum_{j}\frac{\exp(g_{n,j})}{\sum_{m}\exp(g_{n,m})}=1$, we get:
@@
$\left|\frac{\partial\mathsf{J}_{NCE}^{global}}{\partial\theta_{k}}\right|\leq\mathsf{E}_{(x,y)}[B_{g}+B_{g}\cdot 1]=2B_{g}.$
Thus, gradients are bounded.
For the asymmetric NCE loss, $\mathsf{J}_{NCE}=\mathsf{J}^{(+)}+\sum_{i}\mathsf{J}^{(-)}$, with $p=\sigma(g_{\theta}(x,y))$. The partial derivatives are:
$\frac{\partial\mathsf{J}^{(+)}}{\partial\theta_{k}}=\Big{[}$ $-\gamma_{+}(1-p)^{\gamma_{+}}\log(p)$
$+(1-p)^{\gamma_{+}}\Big{]}\frac{\partial g}{\partial\theta_{k}}$
$\frac{\partial\mathsf{J}^{(-)}}{\partial\theta_{k}}=\Big{[}$ $\gamma_{-}p^{\gamma_{-}}\log(1-p)$
$-p^{\gamma_{-}}\Big{]}\frac{\partial g}{\partial\theta_{k}}$
If $g_{\theta}(x,y)$ (and thus $p$) is bounded away from values that would make $\log(p)$ or $\log(1-p)$ excessively large (e.g., $p\in[\epsilon,1-\epsilon]$ for some small $\epsilon>0$), and if $\frac{\partial g}{\partial\theta_{k}}$ is bounded, then the gradients are bounded. This typically holds for well-behaved neural networks. $\square$
Lemma 1 (Lipschitz Property of $g_{\theta}$)
The matching function $g_{\theta}(x,y)$ is Lipschitz-continuous w.r.t. its parameters $\theta$.
Proof Sketch. Given that $g_{\theta}(x,y)$ is a deep neural network, it is a composition of linear layers and standard Lipschitz-continuous activation functions (e.g., ReLU, Tanh, Sigmoid). Therefore, $g_{\theta}(x,y)$ is Lipschitz-continuous assuming bounded weights and inputs. $\square$
Lemma 2 (Noise Distribution Properties)
The random noise distribution $q(y)$ provides negative samples statistically representative of the overall label space.
Proof Sketch. By sampling uniformly from the entire corpus of labels ($q(y)=1/\|L\|$), the noise distribution broadly covers the characteristics of the full label set. This diversity helps prevent the model from exploiting narrow biases in negative sampling and ensures robust learning signals. $\square$
Lemma 3 (Asymptotic Consistency of NCE)
Under standard conditions (e.g., model expressivity, sufficient samples), minimizing the NCE objective yields an asymptotically consistent estimator of the true data distribution or its associated log-ratio.
Proof Sketch. NCE re-casts density estimation as binary classification between true data and noise. The optimal classifier for this task inherently recovers the ratio of the true to noise densities. This consistency property is well-established in NCE literature, indicating that as data size increases, the estimated distribution converges to the true one. $\square$
Lemma 4 (SGD Convergence to Stationary Point)
Given bounded gradients and Lipschitz continuity of the loss function, stochastic gradient descent (SGD) converges to a stationary point (e.g., local minimum) for non-convex NCE objectives.
Proof Sketch. The conditions of bounded gradients and Lipschitz continuity ensure that SGD steps are well-behaved. For non-convex functions (as is typical for deep learning objectives), SGD is guaranteed to converge to a stationary point. The diversified negative sampling strategy can further aid in navigating the loss landscape and potentially finding better local optima. $\square$
Thus, with bounded gradients and Lipschitz continuity, both our $\alpha$-balanced and asymmetric NCE losses exhibit strong convergence properties under SGD, leading to a stationary point in the parameter space.
@@ D.2 Training Procedure and Hyperparameters
D.3 Qualitative Studies
In this section, we provide a series of illustrative examples (see  Tables 10, 11 and 12) to qualitatively showcase the practical efficacy of our methodology in addressing the compound TTP-Mapping task. We relate our results with the established LLMs, such as ChatGPT 4  [^11] While being extensively studied, we opt to exclude its results in our experiments due to the objective prompt-sensitive performance limitations., which serve as a reference to the overall intricacy of this task.
@@ D.3 Qualitative Studies
For the setup of ChatGPT, for each text, we create a prompt in the following format: What MITRE ATT&CK techniques (TTPs) are explicitly and implicitly mentioned in the following text: [..]. In general, the responses provided by Chat-GPT are remarkable and somewhat accurate in certain instances. However, it is evident that the answers primarily consist of high-level information (sometimes hallucinatory), with a lack of granularity that makes it useful, e.g., for precise modeling of the attack steps.
We provide further a full report analysis of a threat report released by Mandiant (see Wayback machine). Each paragraph in the report is processed by our model and finally techniques were assigned to the tactic bins of the MITRE ATT&CK matrix  [^12] https://attack.mitre.org/, based on a simple assignment algorithm, with two constraints (1) maximize total relevant score (of each TTP) in the bins and (2) maximize total number of TTP-occurrences in the bins (i.e., a TTP can occur in more than one paragraph). Further details are in Table 9.
Table 9: A full report analysis of the Mandiant threat report (see Wayback machine). We compare our results with the list of TTPs explicitly provided by the same report, Appendix section, with that, we achieve 90% recall, missing only one technique (Non-Standard Port). All the extracted TTPs from the model are further examined and confirmed correct by our security experts.
Tactics Techniques
Reconnaissance Vulnerability Scanning (T1595.002)
Resource Development Vulnerabilities (T1588.006) Exploits (T1588.005)
Initial Access External Remote Services (T1133) Exploit Public-Facing Application (T1190)
Execution Windows Command Shell (T1059.003) Exploitation for Client Execution (T1203)
Persistence BITS Jobs (T1197) Windows Service (T1543.003)
Privilege Escalation Process Hollowing (T1055.012) Exploitation for Privilege Escalation (T1068)
Defense Evasion Obfuscated Files or Information (T1027) Deobfuscate/Decode Files or Information (T1140)
Discovery System Information Discovery (T1082) File and Directory Discovery (T1083)
Lateral Movement Exploitation of Remote Services (T1210) Lateral Tool Transfer (T1570)
Command and Control Ingress Tool Transfer (T1105) File Transfer Protocols (T1071.002)
Table 10: Example 1: A threat report text is analyzed by ChatGPT and $\alpha$-balanced. Here, an important technique is the Masquerading, described by MITRE as: “Adversaries may attempt to manipulate features of their artifacts to make them appear legitimate or benign to users and/or security tools. Masquerading occurs when the name or location of an object, legitimate or malicious, is manipulated or abused for the sake of evading defenses and observation. This may include manipulating file metadata, tricking users into misidentifying the file type, and giving legitimate task or service names.“ This technique is missed by ChatGPT, identified by our model, and is ranked higher than explicit TTPs i.e., Powershell or Visual Basic.
@@
Text After the .NET PE file has been run, we observed the same behavior as the above QUADAGENT sample of dropping a PowerShell script with the filename SystemDiskClean.ps1 alongside a VBScript file with the same name. The C2 techniques remained identical, with the only change being the server which became cpuproc[.]com.}
ChatGPT inline,color=green!20!whiteinline,color=green!20!whitetodo: inline,color=green!20!whiteBased on the provided text, we can identify several MITRE ATT&CK techniques, both explicitly and implicitly: Execution (T1059): The text mentions the execution of a .NET PE file, indicating the execution technique. Scripting (T1064): The dropping of a PowerShell script (SystemDiskClean.ps1) and a VBScript file suggests the use of scripting techniques. Command and Scripting Interpreter (T1059.003): Specifically, the use of PowerShell in dropping the SystemDiskClean.ps1 file implies the use of command and scripting interpreters. Deobfuscate/Decode Files or Information (T1140): The use of a .NET PE file suggests potential obfuscation or encoding that may need to be deobfuscated or decoded. Command and Control (T1043): The mention of a change in the C2 server (cpuproc[.]com) indicates the use of a different command and control infrastructure.
$\alpha$-balanced inline,color=purple!20!whiteinline,color=purple!20!whitetodo: inline,color=purple!20!whiteThe top-n ranking is: 1. Masquerading (T1036), score: 0.68 2. PowerShell (T1059.001), score: 0.58 3. Visual Basic (T1059.005), score: 0.52 4. Obfuscated Files or Information (T1027), score: 0.47
Table 11: Example 2: A threat report text is analyzed by ChatGPT and $\alpha$-balanced. Here, ChatGPT provides a correct justification on the use of the common HTTP port 443. This is a false positive from our model, as a simple text matching fails to reason whether a port number is common or uncommon. Our model however, is still able to identify the important technique Ingress Tool Transfer, which described by MITRE as: “Adversaries may transfer tools or other files from an external system into a compromised environment. Tools or files may be copied from an external adversary-controlled system to the victim network through the command and control channel or through alternate protocols such as ftp. Once present, adversaries may also transfer/spread tools between victim devices within a compromised environment.
Text The Meterpreter downloader ‘TzGG’ was configured to communicate with 91.208.184[.]78 over port 443 to download the shellcode (MD5: 659bd19b562059f3f0cc978e15624fd9) for Cobalt Strike BEACON (trial-version).
ChatGPT inline,color=green!20!whiteinline,color=green!20!whitetodo: inline,color=green!20!whiteBased on the provided text, we can identify several MITRE ATT&CK techniques, both explicitly and implicitly: Data Obfuscation (T1001): The use of the term ”shellcode” implies the presence of binary code that is typically obfuscated and designed to be directly executed. Standard Application Layer Protocol (T1071): The communication with IP address 91.208.184[.]78 over port 443 suggests the use of HTTPS, which is a standard application layer protocol. Command and Control (T1043): The Meterpreter downloader ’TzGG’ communicating with an external server over port 443 indicates a command and control channel. Connection Proxy (T1090.002): The use of port 443 may involve the exploitation of the encrypted communication channel often used for securing web traffic, serving as a form of connection proxy. Impersonation of Legitimate Entity (T1090.002): The mention of Cobalt Strike BEACON (trial-version) suggests the potential for attackers to impersonate legitimate entities or tools. Application Layer Protocol (T1071.001): The use of port 443 for communication is specifically associated with the HTTPS protocol at the application layer. Download Evil Content (T1105): The Meterpreter downloader ’TzGG’ is configured to download shellcode for Cobalt Strike BEACON, indicating the download of malicious content. Remote Access Tools (T1219): Cobalt Strike BEACON is a well-known remote access tool often used by adversaries for post-exploitation activities.
$\alpha$-balanced inline,color=purple!20!whiteinline,color=purple!20!whitetodo: inline,color=purple!20!whiteThe top-n ranking is: 1. Non-Standard Port (T1105), score: 0.76 2. Ingress Tool Transfer (T1059.001), score: 0.71 3. Application Layer Protocol (T1071), score: 0.59 4. Web Protocols (T1071.001), score: 0.57
Table 12: Example 3: A threat report text is analyzed by ChatGPT and $\alpha$-balanced. Here interestingly, both answers show certain consensus. Our model is still able to match to low-level techniques i.e., Symmetric Cryptography, derived from the use of RC4 for encryption.
Text Unlike previously identified RedDelta PlugX samples, this one uses a legitimate Avast Proxy executable for the second DLL side-loading phase rather than the legitimate Adobe executable used in previously analyzed samples. When loaded into memory, the PlugX payload uses www.systeminfor[.]com for command and control — the same domain used across the Catholic church-themed PlugX samples. Across all four newly analyzed samples, the first stage DLL loader files share an identical, uncommon import hash and rich header with previously observed first stage RedDelta PlugX DLLs. The loaded PlugX payload also matches the custom RedDelta PlugX variant previously described, using RC4 encryption for C2 communications with the same hardcoded RC4 passphrase and configuration block decoding function.
ChatGPT inline,color=green!20!whiteinline,color=green!20!whitetodo: inline,color=green!20!whiteBased on the provided text, we can identify several MITRE ATT&CK techniques, both explicitly and implicitly: DLL Side-Loading (T1073.002): The mention of using a legitimate Avast Proxy executable for the second DLL side-loading phase indicates the use of DLL side-loading, a technique where a malicious DLL is loaded into a legitimate process. Masquerading (T1036): The use of a legitimate Avast Proxy executable for the second DLL side-loading phase is a form of masquerading or disguising the malicious activity to appear legitimate. Command and Control (T1043): The PlugX payload uses the domain www.systeminfor[.]com for command and control, indicating communication with an external server. Standard Application Layer Protocol (T1071): The use of a domain for command and control suggests the use of standard application layer protocols for communication. Data Obfuscation (T1001): The use of RC4 encryption for C2 communications indicates the obfuscation of communication traffic. RC4 (T1027): The mention of using RC4 encryption for C2 communications specifically refers to the use of the RC4 symmetric encryption algorithm. Data Encoding (T1132): The use of RC4 encryption can be considered a form of data encoding for communication between the implant and the C2 server. Encoded/Encrypted Files or Information (T1022): The RC4 encryption is used for encrypting communications between the PlugX payload and the C2 server, securing the information exchanged. Registry Run Keys / Startup Folder (T1060): While not explicitly mentioned in the provided text, the persistence mechanism used by PlugX (loading into memory) often involves leveraging registry run keys or startup folders.
$\alpha$-balanced inline,color=purple!20!whiteinline,color=purple!20!whitetodo: inline,color=purple!20!whiteThe top-n ranking is: 1. DLL Side-Loading (T1574.002), score: 0.81 2. Obfuscated Files or Information (T1027), score: 0.56 3. DLL Search Order Hijacking (T1574.001), score: 0.52 4. Encrypted Channel (T1573), score: 0.49 5. Symmetric Cryptography (T1573.001), score: 045 6. Deobfuscate/Decode Files or Information (T1140), score: 0.32 7. Masquerading (T1036), score: 0.32 8. Registry Run Keys / Startup Folder (T1547.001), score: 0.31